OpenStudy (anonymous):

A film set designer is using white and colored tiles in a pattern to create paths of different lengths. If x is the length of the path in feet, the number of colored tiles needed to make it is calculated using the rule (6x2y3)(4xy3) where y represents the number of white tiles. Which simplified expression represents the number of colored tiles used for a path of length x feet? 24x2y6 24x3y6 24x2y9 24x3y9

OpenStudy (anonymous):

@insanitycat

OpenStudy (anonymous):

(6x2y3)(4xy3) so when we multiply monomials like this, we multiply the coefficients like normal so 6*4=24 but when variables are multiplied, the exponents are added so we have x^2*x^1=x^3 and y^3+y^3=y^6 and now we combine the terms 24x^3y^6 do you understand?

OpenStudy (anonymous):

yes I do thank you so much